# Transformer-based RGB-T Tracking with Channel and Spatial Feature Fusion

### v1

@@ III-C Spatial feature fusion module
Our proposed SFM module aims to model spatial relationships in cross-modal features and to integrate the channel and spatial features of the RGB and TIR modalities through a convolutional feedforward network. As shown in Figure 4, the SFM contains a local perception unit (LPU), a cross-attention module and a convolutional feedforward network (CFN).
@@
Then a cross-attention module is used to achieve spatial modeling of cross-modal features. Specifically, a non-linear full-connection layer is used and represented as $X_{r}^{\text{res}},X_{r}^{\text{attn}}=\text{Split}(\text{Relu}(\text{Linear}^{\text{s}}(X_{r}^{\text{lpu}})))$, where $\text{Linear}^{\text{s}}$ is a linear layer shared by RGB and TIR branches. After that, a linear layer is used to obtain Query: $X_{r}^{q}=\text{Linear}(X_{r}^{\text{attn}})$, another linear layer is used to obtain Key: $X_{r}^{k}$ and Value: $X_{r}^{v}$, represented as $X_{r}^{k},X_{r}^{v}=\text{Split}(\text{Linear}(X_{s}^{\text{attn}})$. The TIR $X_{t}^{q},X_{t}^{k},X_{t}^{v}$ are obtained through the same way.
@@
Then we use residual connections and a shared linear layer to adjust the features and add them up, the process is described as:
@@
Then, we add the features $X_{r}^{\text{add}}$ and $X_{t}^{\text{add}}$ to achieve feature point-level interaction between the multimodal features.
@@
Due to the differences between RGB and TIR features, the directly added feature requires further spatial and channel integration. Therefore, a convolutional feedforward network (CFN) is introduced. In the CFN, we first perform spatial integration on the fused feature. The process can be written as follows:
@@
Then, we nonlinearly integrate the channel feature of the fused feature, represented as:
@@
In addition, we use a 1x1 convolutional block and a skip connection with a 1x1 convolutional block to adjust the features, the process is described as:
@@ IV Experiments
IV-A Datasets and Evaluation Metrics
LasHeR [28] is a large-scale dataset benchmark that contains 1224 sequences with 730K frame pairs. It is divided into 19 challenging attributes, which are No Occlusion (NO), Partial Occlusion (PO), Total Occlusion (TO), Hyaline Occlusion (HO), Out-of-View (OV), Low Illumination (LI), High Illumination (HI), Abrupt Illumination Variation (AIV), Low Resolution (LR), Deformation (DEF), Background Clutter (BC), Similar Appearance (SA), Thermal Crossover (TC), Motion Blur (MB), Camera Moving (CM), Frame Lost (FM), Fast Motion (FM), Scale Variation (SV), Aspect Ratio Change (ARC). RGBT234 [29] is a classic benchmark for RGB-T tracking that contains 234 sequences. RGBT234 [29] has 12 challenging attributes, they are no occlusion (NO), partial occlusion (PO), heavy occlusion (HO), low illumination (LI), low resolution (LR), thermal crossover (TC), deformation (DEF), fast motion (FM), scale variation (SV), motion blur (MB), camera moving (CM) and background clutter (BC). RGBT210 [30] is another classic RGB-T tracking benchmark that contains 210 sequences.
Following the metrics used by current most RGB-T trackers, precision rate (PR) and success rate (SR) are employed to evaluate different trackers on RGBT210 [30] and RGBT234 [29], and precision rate (PR), normalized precision rate (NPR) and success rate (SR) are employed to evaluate the trackers on LasHeR [28].
TABLE I: Comparison with state-of-the-art trackers on RGBT210, RGBT234 and LasHeR [28] testing set. Higher value Indicate better performance. The best two results are shown in red and blue fonts.
@@ IV-A Datasets and Evaluation Metrics
IV-B Implementation Details
CSTNet is implemented based on Ubuntu 20.04, using Python 3.9 and Pytorch 1.13.0. It is trained on a platform with two NVIDIA RTX A6000 GPUs over 20 epochs, which took roughly 6 hours. LasHeR [28] training set is used for training our model. In each epoch, we sample 60k image pairs, with a total batch-size of 64. The total learning rate is set to $2e^{-5}$ for the backbone and $2e^{-6}$ for feature fusion modules. It is decayed by 10× after 15 epochs. We use AdamW as the optimizer with $1e^{-4}$ weight decay. The size of template and search area is set to 128×128 and 256×256, respectively. Our CFM and SFM modules are cascaded and inserted into the 4-th, 7-th, and 10-th layers of the ViT. We use the ViT model weight of the RGB-T tracker [2] as the pre-training weight. CSTNet achieves speed at 62 FPS on GPU GTX 3090Ti.
@@ IV-C1 Comparison methods
To evaluate the effectiveness of our method, we compared it with RGB-T trackers from recent years, including 5 MDNet-based trackers, including TFNet [31], DMCNet [32], APFNet [8], X-Net [33], CAT++ [20], 3 Siamese-based trackers, including SiamCDA [11], SiamMALL [13], DFAT [34], 4 DCF-based trackers MFNet [35], LSAR [36], CMD [37], AMNet [38] and 11 transformer-based trackers, including ProTrack [14], ViPT [1], MACFT [19], RSFNet [39], DFMTNet [40], TBSI [2], STMT [41], BAT [15], TATrack [16], TransAM [38].
@@ IV-C2 Evaluation on LasHeR Dataset
The trackers are evaluated on LasHeR [28] and the results are reported on Table I. CSTNet achieves state-of-the-art performance. Compare to other Transformer-based trackers, CSTNet outperforms TATrack [16] and TransAM [38] by both 1.3% in PR, 1.2% and 1.9% in NPR, 1.1% and 1.3% in SR, respectively. In addition, the PR and SR of CSTNet are 1.3% and 0.9% higher than BAT [15]. In addition, CSTNet surpasses all other MDNet-based and Siamese-based methods. The results show that CSTNet achieves better cross-modal feature interaction and better performance.
In addition, we evaluate CSTNet, TBSI [2] and BAT [15] on 19 attributes of LasHeR [28] dataset. As shown in Figure 5, CSTNet outperforms other transformer-based trackers in many attributions. Compared to TBSI [2], CSTNet achieves more competitive performance in challenges such as occlusion (HO, PO and SO), thermal crossover (TC), out-of-view (OV) and motion blur (MB). Compare to BAT [15], CSTNet has better performance in attributes such as heavy occlusion (HO), abrupt illumination variation (AIV), and background clutter (BC). However, the PR and SR of CSTNet on ARV, FL, SA, AIV and LI attributes are insufficient, which proves that CSTNet needs to further improve the robustness and feature discrimination of the model.
TABLE II: Ablation studies of our proposed CFM and SFM modules in the tracking pipeline.
@@
TABLE III: Ablation of the different methods for fusing search area cross-modal features.
@@
TABLE IV: Ablation of the different components of CFM.
@@
TABLE V: Ablation of the different components of SFM.
@@
IV-C3 Evaluation on RGBT234 Dataset
The trackers are evaluated on RGBT234 [29] and the results are reported on Table I. Compared to the TransAM [38], CSTNet improves the PR by 0.7%. Compared to the BAT [15], CSTNet improves the PR and SR by 1.6% and 1.1%, respectively. The results show that CSTNet sets a new state-of-the-art benchmark.
We further evaluate our proposed CSTNet, TBSI [2], and mfDiMP [42] on 12 attributes of RGBT234 [29] dataset. As shown in Figure 6, CSTNet achieves better performance on most attributes compared to other trackers. However, it has insufficient success rate in heavy occlusion (HO) and low resolution (LR) background clutter (BC) and fast motion (FM). This indicates that it has shortcomings in dealing with the challenges such as missing key features and interference from background features.
IV-C4 Evaluation on RGBT210 Dataset
The trackers are evaluated on RGBT210 and the results are reported on Table I. CSTNet outperforms all other RGB-T trackers. The PR, SR of CSTNet are 0.7%, 1.0% higher than TBSI [2]. In addition, CSTNet achieves an SR score of 63.5%, outperforms STMT [41] and TATrack [16] by 4.0% and 1.7%, respectively. The experimental results demonstrate that CSTNet outperforms all other trackers and achieves state-of-the-art performance.
IV-D Ablation Studies
IV-D1 Component Analysis
We first evaluate the effect of summing RGB and TIR features and the CFM and SFM modules on the performance of the tracker. RGBT denotes adding the RGB and TIR features of the search area output by the shared ViT backbone, and then predicting the target state. The pretrain-model of the baseline is [2]. The impact of different pre-trained models on the tracker will be discussed in Section IV-D4. As shown in Table II, an improvement of 0.5% and 1.0% in term of PR, 0.7% and 0.9% in term of NPR and 0.5% and 0.7% in term of SR is obtained by the contribution of the CFM module and the SFM module, respectively. In addition, they achieved a total of 1.5%, 1.6%, and 1.2% improvement in term of PR, NPR and SR, respectively. Overall, both our proposed CFM and SFM modules are important.
In addition, we separately investigate the impact of concatenating or adding the RGB and TIR features of the search area output from the backbone on the tracking performance. concat is used by TBSI [2], which means that the RGB and TIR features of the search area are concatenated and then a 1x1 convolution is used to reduce the dimension. As shown in Table III, the addition of RGB and TIR features in the search area achieves better performance. Our method avoids the feature representation bottleneck caused by the 1x1 convolution dimension reduction, allowing the model to benefit from the direct addition of RGB and TIR features in the search area.
IV-D2 Ablation of CFM
We evaluate the contributions of the each component of the CFM. As shown in Table IV, When the CFM module is removed from the model, the PR and SR are 69.5% and 55.8%, respectively. Adding the SE module improves the PR by 1.0% and the SR by 0.5%. Then, using the LSA module achieves better performance, with increases of 0.3% for SR and 0.4% for PR, respectively. Next, using the GIM module to integrate global features results in an improvement of 0.7% for PR and 0.5% for SR, respectively. Overall, all three modules in CFM contribute to improving the performance of our model.
IV-D3 Ablation of SFM
The contribution of each component to the SFM is measured. Table V shows the results. The contribution of the LPU is an improvement of 0.1% for PR. The Cross-attention module improves the PR by 0.4% and SR by 0.3%, respectively. The CFN module improves the PR by 0.5% and SR by 0.4%. Overall, all three components of the SFM are effective.
IV-D4 Ablation of pre-trained models
We explore the impact of different pre-trained models on our proposed CSTNet in terms of performance. CSTNet without pretraining and CSTNet with ImageNet pretraining are unable to converge in 20 epochs, therefore we increase their learning rate by 10x. The results are shown in Table VI, CSTNet with RGB-T pretraining achieves the best performance. It outperforms CSTNet with SOT pretraining by 1.8% for PR and 1.4% for SR. This shows that using the RGB-T pre-trained model can further improve the performance of the tracker.
TABLE VI: Ablation studies of different pre-trained models.
Pre-train Model P NP SR
- 55.6 51.7 42.2
ImageNet [43] 59.1 56.0 47.3
SOT [22] 69.7 65.9 55.8
RGBT [2] 71.5 67.9 57.2
IV-D5 Inserting Layers of CFM and SFM
In addition, We conduct ablation experiments on different insert layers of our proposed CFM and SFM modules. The results are reported in Table VII. Inserting the CFM and SFM modules in the 4-th layer of ViT improves the PR by 0.4% and the SR by 0.3%. Inserting the modules in the 7th and 10th layers of ViT further improves the performance of the tracker. Overall, inserting our CFM and SFM modules into these three layers effectively achieves feature fusion at different stages and improves model performance.
TABLE VII: Ablation studies of different insert layers of our CFM and SFM modules.
Layers P NP SR
@@ IV-D5 Inserting Layers of CFM and SFM
IV-E Visualization
IV-E1 Heatmap

### v2

@@ III-C Spatial feature fusion module
Then a cross-attention module is used to achieve spatial modeling of cross-modal features. Specifically, a non-linear full-connection layer is used and represented as $X_{r}^{\text{res}},X_{r}^{\text{attn}}=\text{Split}(\text{Relu}(\text{Linear}^{\text{s}}(X_{r}^{\text{lpu}})))$, where $\text{Linear}^{\text{r}}$ is a linear layer shared by RGB and TIR branches. After that, a linear layer is used to obtain Query: $X_{r}^{q}=\text{Linear}(X_{r}^{\text{attn}})$, another linear layer is used to obtain Key: $X_{r}^{k}$ and Value: $X_{r}^{v}$, represented as $X_{r}^{k},X_{r}^{v}=\text{Split}(\text{Linear}(X_{r}^{\text{attn}})$. The TIR $X_{t}^{q},X_{t}^{k},X_{t}^{v}$ are obtained through the same way.
@@
Then we use residual connections and a shared linear layer to adjust the features and sum the features. The reason for using sum instead of concat is to maintain channel consistency between the semantic structure of the RGB features and the TIR features of the target. The shared layer improves the learning of unified features in cross-modal features of the target. The process is described as:
@@
For the same reasons as above, we add the features $X_{r}^{\text{add}}$ and $X_{t}^{\text{add}}$ to achieve feature point-level interaction between the multimodal features.
@@
Due to the differences between RGB and TIR features, the directly added feature requires further spatial and channel integration. Therefore, a convolutional feedforward network (CFN) is introduced. The CFN improves the spatial and channel representation of key features by learning spatial and channel decoupling of cross-modal features. In the CFN, we first perform spatial integration on the merged feature. The process can be written as follows:
@@
Then, we nonlinearly integrate the channel feature of the merged feature, represented as:
@@
In addition, we use a 1x1 convolutional block and a skip connection with a 1x1 convolutional block to adjust the features, the process is described as follows:
@@ IV Experiments
TABLE I: Comparison with state-of-the-art trackers on RGBT210 [29], RGBT234 [30] and LasHeR [31] testing set. Higher value Indicate better performance. The best two results are shown in red and blue fonts.
@@
IV-A Datasets and Evaluation Metrics
LasHeR [31] is a large-scale dataset benchmark that contains 1224 sequences with 730K frame pairs. It is divided into 19 challenging attributes, which are No Occlusion (NO), Partial Occlusion (PO), Total Occlusion (TO), Hyaline Occlusion (HO), Out-of-View (OV), Low Illumination (LI), High Illumination (HI), Abrupt Illumination Variation (AIV), Low Resolution (LR), Deformation (DEF), Background Clutter (BC), Similar Appearance (SA), Thermal Crossover (TC), Motion Blur (MB), Camera Moving (CM), Frame Lost (FM), Fast Motion (FM), Scale Variation (SV), Aspect Ratio Change (ARC). RGBT234 [30] is a classic benchmark for RGB-T tracking that contains 234 sequences. RGBT234 [30] has 12 challenging attributes, they are no occlusion (NO), partial occlusion (PO), heavy occlusion (HO), low illumination (LI), low resolution (LR), thermal crossover (TC), deformation (DEF), fast motion (FM), scale variation (SV), motion blur (MB), camera moving (CM), and background clutter (BC). RGBT210 [29] is another classic RGB-T tracking benchmark that contains 210 sequences. VTUAV [42] is a large-scale RGB-T tracking dataset in aerial view and its short-term tracking test subset contains 176 sequences. GTOT [43] is a classic RGB-T tracking benchmark consisting of 50 short-term sequences.
Following the metrics used by current most RGB-T trackers, precision rate (PR) and success rate (SR) are employed to evaluate different trackers on RGBT210 [29], RGBT234 [30], and GTOT [43], and precision rate (PR), normalized precision rate (NPR) and success rate (SR) are employed to evaluate the trackers on LasHeR [31], and maximum precision rate (MPR) and maximum success rate (MSR) are employed to evaluate the trackers on VTUAV [42].
@@ IV-A Datasets and Evaluation Metrics
IV-B Implementation Details
CSTNet is implemented on Ubuntu 20.04 using Python 3.9 and Pytorch 1.13.0. It is trained on a platform with two NVIDIA RTX A6000 GPUs over 20 epochs, which took roughly 6 hours. The LasHeR [31] training set is used to train our model. In each epoch, we sample 60k image pairs, with a total batch-size of 64. The total learning rate is set to $2e^{-5}$ for the backbone and $2e^{-6}$ for feature fusion modules. It decays 10× after 15 epochs. We use AdamW as the optimizer with $1e^{-4}$ weight decay. The sizes of template and search area are set to 128×128 and 256×256, respectively. Our CFM and SFM modules are cascaded and inserted into the 4-th, 7-th, and 10-th layers of the ViT.
CSTNet-small removes the CFM and SFM modules from the network and uses the CSTNet weights as the pre-training weights, while its training settings are the same as CSTNet.
@@ IV-C1 Comparison methods
To evaluate the effectiveness of our method, we compare CSTNet and CSTNet-small with RGB-T trackers from recent years, including 5 MDNet-based trackers, including TFNet [32], DMCNet [33], APFNet [9], X-Net [34], CAT++ [21], 3 Siamese-based trackers, including SiamCDA [12], SiamMALL [14], DFAT [35], 4 DCF-based trackers MFNet [36], LSAR [37], CMD [38], AMNet [6], and 11 transformer-based trackers, including ProTrack [15], ViPT [1], MACFT [20], RSFNet [39], DFMTNet [40], TBSI [2], STMT [41], BAT [16], TATrack [17], TransAM [6].
TABLE II: Comparison with state-of-the-art trackers on GTOT [43]. The best two results are shown in red and blue fonts.
Baseline Backbone GTOT
PR SR
TFNet [32] MDNet VGG-M 72.4 88.6
APFNet [9] MDNet VGG-M 73.7 90.5
DMCNet [33] MDNet VGG-M 73.8 90.9
SiamCDA [12] Siamese Res-50 87.7 73.2
DFAT [35] Siamese Res-50 89.3 72.3
SiamMLAA [14] Siamese Res-50 91.3 75.1
LSAR [37] DCF VGG-16 70.3 85.5
CMD [38] DCF Res-18 89.2 73.4
AMNet [6] DCF Res-50 91.1 76.1
RSFNet [39] Transformer ViT-B 92.1 75.3
TransAM [6] Transformer ViT-B 92.9 77.4
CSTNet-small Transformer ViT-B 91.5 75.4
CSTNet Transformer ViT-B 93.0 76.7
TABLE III: Comparison of parameters, Flops, and FPS on GPU GTX 3090Ti and GPU GTX 2080.
3090Ti / 2080 Params Flops
torch onnxruntime
TBSI [2] 60 / 28 71 / 40 201.98M 82.5G
CSTNet 62 / 36 74 / 44 144.00M 74.7G
CSTNet-small 94 / 47 126 / 65 92.12M 56.4G
Figure 7: Comparison with state-of-the-art trackers on VTUAV [42]. The trackers include HMFT [42], Stark [45], TransT [46], mfDiMP [44], FSRPN [47], TCTrack [48], ADRNet [49], DAFNet [7].
@@ IV-C2 Evaluation on LasHeR Dataset
The trackers are evaluated on LasHeR [31] and the results are reported in Table I. CSTNet achieves state-of-the-art performance. Compared to other Transformer-based trackers, CSTNet outperforms TATrack [17] and TransAM [6] by 1.3% in PR, 1.2% and 1.9% in NPR, and 1.1% and 1.3% in SR, respectively. In addition, CSTNet’s PR and SR are 1.3% and 0.9% higher than BAT [16]. In addition, CSTNet surpasses all other MDNet-based and Siamese-based methods. CSTNet-small achieves performance close to that of BAT. The results show that CSTNet achieves better cross-modal feature interaction and better performance.
In addition, we evaluate CSTNet, TBSI [2], and BAT [16] on 19 attributes of the LasHeR [31] dataset. As shown in Figure 5, CSTNet outperforms other transformer-based trackers on many attributes. Compared to TBSI [2], CSTNet achieves more competitive performance in challenges such as occlusion (HO, PO, and SO), thermal crossover (TC), out-of-view (OV) and motion blur (MB). Compared to BAT [16], CSTNet has better performance in attributes such as heavy occlusion (HO), abrupt illumination variation (AIV), and background clutter (BC). However, the PR and SR of CSTNet on ARV, FL, SA, AIV, and LI attributes are insufficient, which proves that CSTNet needs to further improve the robustness and feature discrimination of the model.
IV-C3 Evaluation on RGBT234 Dataset
The trackers are evaluated on RGBT234 [30] and the results are reported in Table I. Compared to the TransAM [6], CSTNet improves the PR by 0.7%. Compared to the BAT [16], CSTNet improves the PR and SR by 1.6% and 1.1%, respectively. CSTNet outperforms CSTNet-small by 1.7% in PR and 1.1% in SR. The results show that CSTNet sets a new state-of-the-art benchmark.
We further evaluate our proposed CSTNet, TBSI [2], and mfDiMP [44] on 12 attributes of the RGBT234 [30] dataset. As shown in Figure 6, CSTNet achieves better performance on most attributes compared to other trackers. However, it has an insufficient success rate in heavy occlusion (HO), and low resolution (LR) background clutter (BC), and fast motion (FM). This indicates that it has shortcomings in dealing with challenges such as missing key features and interference from background features.
IV-C4 Evaluation on RGBT210 Dataset
The trackers are evaluated on RGBT210 [29] and the results are reported in Table I. CSTNet outperforms all other RGB-T trackers. The PR, SR of CSTNet are 0.7%, 1.0% higher than TBSI [2]. In addition, CSTNet achieves an SR of 63.5%, outperforming STMT [41] and TATrack [17] by 4.0% and 1.7%, respectively. The experimental results demonstrate that CSTNet outperforms all other trackers and achieves state-of-the-art performance.
IV-C5 Evaluation on VTUAV-ST Dataset
The trackers are evaluated on VTUAV [42] and the results are reported in Table 7. CSTNet and CSTNet-small outperform all other RGB-T trackers. Compared to HMFT [42], CSTNet improves the MPR by 1.7% and the MSR by 3.3%. Compared to CSTNet, CSTNet-small improves the MPR by 4.7% and the MSR by 3.3%. The experimental results demonstrate that CSTNet-small outperforms all other trackers and achieves state-of-the-art performance.
IV-C6 Evaluation on GTOT Dataset
The trackers are evaluated on GTOT [43] and the results are reported in Table II. CSTNet outperforms all other RGB-T trackers in PR. Compared to RSFNet [39], CSTNet improves the PR and SR by 0.9% and 1.4%, respectively. The experimental results demonstrate that CSTNet and CSTNet-small achieves state-of-the-art performance.
TABLE IV: Ablation studies of our proposed CFM and SFM modules in the tracking pipeline.
@@ IV-C6 Evaluation on GTOT Dataset
TABLE V: Ablation of the different methods for fusing search area cross-modal features.
@@
TABLE VI: Ablation of the different components of CFM.
@@
TABLE VII: Ablation of the different components of SFM.
@@
TABLE VIII: Ablation studies of different insert layers of our CFM and SFM modules.
Layers PR NPR SR
@@
TABLE IX: Ablation studies of different pre-trained models.
Pre-train Model PR NP SR
- 55.6 51.7 42.2
ImageNet [50] 59.1 56.0 47.3
SOT [23] 69.7 65.9 55.8
RGBT [2] 71.5 67.9 57.2
IV-D Parameters and Flops
We also evaluate the Parameters, Flops, and FPS of the models. The results are shown in Table III. Compared to our baseline [2], our CSTNet and CSTNet-small achieve better performance, reducing approximately 25% and 50% of the parameters and 10% and 30% of the Flops, respectively. Our CSTNet and CSTNet-small achieve 62 FPS and 94 FPS, respectively, on the GPU GTX 3090Ti. We use onnxruntime to further accelerate the model. The accelerated CSTNet-small achieves a speed of 126 FPS. We also test the speed of the trackers on the GPU GTX2080. CSTNet and CSTNet-small run at a speed of 36 and 47 FPS respectively. After the deployment, their speeds reached 44 FPS and 65 FPS respectively. In addition, we deploy our CSTNet and CSTNet-small on NVIDIA Jetson Xavier. CSTNet and CSTNet-small run at speeds of 21 FPS and 33 FPS, demonstrating the feasibility of our method in practical applications.
IV-E Ablation Studies
IV-E1 Component Analysis
First, we evaluate the effect of summing the RGB and TIR features and the CFM and SFM modules on the performance of the tracker. RGBT denotes adding the RGB and TIR features of the search area output by the shared ViT backbone, and then predicting the target state. The pretrain-model of the baseline is [2]. As shown in Table IV, an improvement of 0.5% and 1.0% in terms of PR, 0.7% and 0.9% in terms of NPR, and 0.5% and 0.7% in terms of SR is obtained by the contribution of the CFM module and the SFM module, respectively. In addition, they achieve a total improvement of 1.5%, 1.6%, and 1.2% in terms of PR, NPR, and SR, respectively. In addition, the 1-2% performance degradation of CSTNet when using only the SFM module confirms the discussion in TBSI that cross-attention may not be suitable for RGB-T tracking. In fact, the cross-modal features integrated through CFM can lead to positive gains with the cross-attention based SFM module. This further illustrates the importance of cascading CFM and SFM modules. They are interdependent, rather than simply cascading two modules to improve features separately, which distinguishes our method from other cross-attention-based methods. Overall, both our proposed CFM and SFM modules are important.
In addition, we separately investigate the impact of concatenating or adding the RGB and TIR features of the search area output from the backbone on the tracking performance. concat is used by TBSI [2], which means that the RGB and TIR features of the search area are concatenated and then a 1x1 convolution is used to reduce the dimension. As shown in Table V, the addition of RGB and TIR features in the search area achieves better performance. Our method avoids the feature representation bottleneck caused by the 1x1 convolution dimension reduction, allowing the model to benefit from the direct addition of RGB and TIR features in the search area.
IV-E2 Ablation of CFM
We evaluate the contributions of each component of the CFM. As shown in Table VI, when the CFM module is removed from the model, the PR and SR are 69.5% and 55.8%, respectively. Adding the SE module improves the PR by 1.0% and the SR by 0.5%. Then, using the LSA module improves performance by 0.3% for SR and 0.4% for PR. Next, using the GIM module to integrate global features results in an improvement of 0.7% for PR and 0.5% for SR. Overall, all three modules in CFM contribute to improving the performance of our model.
IV-E3 Ablation of SFM
The contribution of each component to the SFM is measured. Table VII shows the results. The contribution of the LPU is a 0.1% improvement in PR. The cross-attention module improves the PR by 0.4% and SR by 0.3%. The CFN module improves the PR by 0.5% and SR by 0.4%. Overall, all three components of the SFM are effective.